# Foundation Models for Rapid Autonomy Validation

### v1

@@ Abstract
We are motivated by the problem of autonomous vehicle performance validation. A key challenge is that an autonomous vehicle requires testing in every kind of driving scenario it could encounter, including rare events, to provide a strong case for safety and show there is no edge-case pathological behavior. Autonomous vehicle companies rely on potentially millions of miles driven in realistic simulation to expose the driving stack to enough miles to estimate rates and severity of collisions. To address scalability and coverage, we propose the use of a behavior foundation model, specifically a masked autoencoder (MAE), trained to reconstruct driving scenarios. We leverage the foundation model in two complementary ways: we (i) use the learned embedding space to group qualitatively similar scenarios together and (ii) fine-tune the model to label scenario difficulty based on the likelihood of a collision upon re-simulation. We use the difficulty scoring as importance weighting for the groups of scenarios. The result is an approach which can more rapidly estimate the rates and severity of collisions by prioritizing hard scenarios while ensuring exposure to every kind of driving scenario.
@@ I Introduction
As human operators are removed from controlling autonomous vehicles and driving software matures to carry passengers, simulated behavior validation takes an ever more central role. Behavior validation entails justifying that the software stack responsible for executive control of the vehicle can achieve a desired performance target across the expected exposure of an operating design domain (ODD). Without a human operator either physically present or remotely assisting, the validation stack must both exercise all expected behaviors and return a set of concrete metrics in a trustworthy way.
@@
In this work, we focus on the setting where a developer has collected significant driving logs but seeks to prioritize their simulation. This prioritization may be used within a fixed simulation budget to maximize effective validation, to reduce a set of simulations in order to save compute spend or total validation latency, or a combination thereof. While prior works have performed searches to find adversarial conditions that stress the autonomy stack [1, 2, 3, 4], choosing from a fixed set of logs allows for estimating performance on the distribution of expected scenarios, as they come from real driving scenes.
Figure 1: Overview of the proposed validation process. We embed a scenario using a pretrained encoder $\mathcal{E}$. That embedding is concatenated with a difficulty score from a predictor head $\mathcal{D}$ taking in the embedding as an input. For a scenario set, these concatenated embeddings are clustered by an algorithm $\mathcal{C}$ and finally sampled by a sampler $\mathcal{S}$ at validation time. A metrics stage $\mathcal{M}$ simulates and computes metrics over the sampled scenarios.
@@
  • The sampling procedure should return the exposure to each scenario type to calculate a weighted validation metric.
@@ II Related Work
AV Validation & Verification. ISO 26262 is a functional safety standard for automotive electronic and electrical systems, focusing on lifecycle management and risk assessment. ISO 26262 sets the limit for the acceptable ratio of faults as 10 Failures in Time (FIT), meaning that given $10^{9}$ hours of operation no more than 10 faults should be observed. However, AVs are also generally unsafe to test on the road at that scale and modern AVs undergo extremely dynamic and fast-paced engineering cycles that necessitate frequent re-verification and validation [68, 69]. Simulation has emerged as a partial solution for AV V&V that is rapid, scalable, and safe compared to real-world testing. Plenty of open-source simulators exist for AV testing [70, 71, 72, 73, 74] with some including benchmarks for autonomous driving quality [75, 76]. Despite being more scalable than real-world testing, simulated driving can still be prohibitively expensive at scale. A standard approach is to mine for difficult scenarios from existing driving logs [1, 2, 3, 4], or synthesize difficult scenarios often using ML [77, 78, 79, 80, 81, 82, 83] with [84] showing off a controllable diffusion-based approach allowing users to specify desired scenario properties. Fault injection is another common approach for creating difficult test scenarios, with frameworks such as [85, 86, 87, 88, 89, 90] using algorithms like Bayesian Optimization to rapidly hunt for faults that lead to system failures. Other approaches specifically target ML-based components by inserting sensor noise [91, 92], with “white-box” frameworks like [93, 94] adversarially mutating inputs in an attempt to maximize “neuron coverage” in the system to be tested. Formal verification methods typically attempt to give rigorous guarantees on system safety by using mathematical tools like Lyapunov functions [95, 96, 97] and Satisfiability Modulo Theories [98, 99], or logical modelling and proofs [100, 101, 102, 103, 104].
@@ III-A1 Tracks
The track tensor $X_{T}\in\mathbb{R}^{N_{T}\times T\times D_{T}}$ contains all tracks from the perception system, including classes such as vehicles, pedestrians, cones, lane dividers, and others. For each track-time step, we encode the pose, velocity, acceleration, extents, and class into a single $D_{T}$-dimensional vector. We assign each unique track to a row in $X_{T}$. If there are fewer than $N_{T}$ tracks or a track is not visible at all time steps, $X_{T}$ is zero-padded where necessary.
@@ III-A2 Traffic Signals
The signal tensor $X_{S}\in\mathbb{R}^{N_{S}\times T\times D_{S}}$ contains rows where each signal-time step is a $D_{S}$-vector encoding pose and perception label for a single signal. Since we use a map with signal locations, no time steps are padded as they can be set with the pose and an UNKNOWN label.
@@ III-B Time Invariant Inputs
To encode the road network, we adopt an approach similar to [5]. We extract vectorized polylines representing the centers of lanes and annotation marks such as stop lines, parking spaces, and crosswalks. The polylines are represented in three tensors:
  • Coordinate frames for each polyline $X_{F}\in\mathbb{R}^{N_{Z}\times D_{F}}$ containing $x,y$ locations and sine and cosine of the rotation to the frame
  • Class labels $X_{L}\in\mathbb{R}^{N_{Z}\times D_{L}}$
  • Sets of polyline points $X_{P}\in\mathbb{R}^{N_{Z}\times S_{Z}\times D_{P}}$ expressed in the associated coordinate frame with $x,y$ positions, width, and existence
@@ III-C Masking
As is typical in an MAE setup, we randomly sample inputs to mask with masking ratio $r$. Masking corresponds to replacing the encoded features’ vectors with a 0-vector. For the time variant masks $M_{T}\in\mathbb{R}^{N_{T}\times T}$ and $M_{S}\in\mathbb{R}^{N_{T}\times T}$, we sample across the first two dimensions of the input independently. For the time invariant mask $M_{Z}\in\mathbb{R}^{N_{Z}}$, we randomly mask over the polylines. If a polyline is masked, the coordinate frame is left unmasked but the labels and points are all masked, including the existence.
@@ III-D Model
  • The inputs are not tokenized and of varying sizes, so we project them into a shared space and back out to their original shape
@@ III-E Loss
We independently sample a mask with ratio $r_{\text{loss}}$ following the same sampling methodology as described previously in the masking section. For all inputs that are covered by this mask, we compute a reconstruction loss. This means that the loss applies to inputs that are masked as well as not masked. The loss is a weighted sum over individual input types:
@@
In addition to the track, signal, and road network losses, we introduce a loss term that focuses on reconstructing ego, as many tasks require a strong representation of ego. Each loss can be decomposed into an L1 loss on the $J$ continuous inputs, such as track pose or polyline point width, and a cross-entropy loss on the $M$ categorical inputs, such as track type:
@@
For the models used in future sections, we set all $\lambda=1$.
@@ IV-A Data
The data for the fine-tuning is similar to data used to train the difficulty model presented in [109]. We use a set of recorded driving logs and their results after resimulation. For a given scenario, we apply a label of 1 if resimulation resulted in a collision and 0 if it did not. In general, it is possible for a driving log to have a simulation result on many different software versions. If so, we add it to the dataset as multiple examples. Note that each would share identical inputs, as we encode the original recorded driving log. For scenarios that do not unimodally fail or succeed, the optimal score is then some real number between 0 and 1.
@@ IV-B Fine-tuning
While there are examples of log divergence and other factors leading to false positive collisions in our training set, the simulator is broadly high quality enough to lead to many simulated collisions being of real concern.
@@ V-A Clustering scenarios
During validation, it is crucial to ensure that every kind of scenario, especially those less common or more complex, is tested to validate changes in behavior comprehensively. After clustering scenarios, we can exposing the autonomy stack to a representative sample from each cluster to cover each of the various behaviors on which validation is required. This intuition motivates a sampling technique where we uniformly sample scenarios across clusters (see Algorithm 1).
@@ V-B Importance sampling
We incorporate a supervised signal, scenario difficulty as discussed in Section IV, to bias sampling towards more useful scenarios. We propose appending (and weighting) the difficulty score as part of the scenario embedded space. Thus clustering will separate based on scenario similarity and difficulty.
With clusters separated in terms of difficulty, we can score each (i.e. according to an average, or even a particular percentile of the difficulty scores within the cluster). The score $w_{i}$ for cluster $c_{i}$ can then be treated as an importance weight upon sampling. The resulting sampling scheme, DICE, presented in Algorithm 2, ensures that the autonomy stack is more frequently exposed to challenging scenarios, which are crucial for thorough validation.
@@ VII Conclusions and Limitations
To illustrate how expensive validation can be, consider a hypothetical example. An AV developer wishes to validate that their driving stack violates a safety metric (measured in miles per violation) no more than once per N miles across geosites using a real-time simulator. If all roads have a 35 mph speed limit and the developer wants to simulate 10 times the desired metric for confidence, this requires $10N/35\approx 0.286N$ driving hours of simulation per geofence. At time of writing, the on-demand cost for an AWS g5.4xlarge node is \$$\$1.624$ per hour [111] and we assume one node fits one concurrent simulation. For each driving stack change, this yields a cost of \$$\$1.858N$. For a violation that should occur fewer than once every 50,000 miles, this would cost roughly \$$\$93,000$ for each code change.
There are many nuances to effective large scale simulation that are not included here, such as running multiple simulations per node, faster than real time simulation, node startup and shutdown costs, data transmission costs, limitations on pipeline width, and simulation fidelity.

### v2

@@ Abstract
We are motivated by the problem of autonomous vehicle performance validation. A key challenge is that an autonomous vehicle requires testing in every kind of driving scenario it could encounter, including rare events, to provide a strong case for safety and show there is no edge-case pathological behavior. Autonomous vehicle companies rely on potentially millions of miles driven in realistic simulation to expose the driving stack to enough miles to estimate rates and severity of collisions. To address scalability and coverage, we propose the use of a behavior foundation model, specifically a masked autoencoder (MAE), trained to reconstruct driving scenarios. We leverage the foundation model in two complementary ways: we (i) use the learned embedding space to group qualitatively similar scenarios together and (ii) fine-tune the model to label scenario difficulty based on the likelihood of a collision upon simulation. We use the difficulty scoring as importance weighting for the groups of scenarios. The result is an approach which can more rapidly estimate the rates and severity of collisions by prioritizing hard scenarios while ensuring exposure to every kind of driving scenario.
@@ I Introduction
As human operators are removed from autonomous vehicles and driving software matures to carry passengers, simulated behavior validation takes an ever more central role. Behavior validation entails justifying that the software stack responsible for executive control of the vehicle can achieve a desired performance target across the expected exposure of an operating design domain (ODD). Without a human operator either physically present or remotely assisting, the validation stack must both exercise all expected behaviors and return a set of concrete metrics in a trustworthy way.
@@
In this work, we focus on the setting where a developer has collected significant driving logs but seeks to prioritize their simulation. This prioritization may be used within a fixed simulation budget to maximize effective validation, to reduce a set of simulations in order to save compute spend or total validation latency, or a combination thereof. Prior works have performed searches to find adversarial conditions that stress the autonomy stack [1, 2, 3, 4]. Our approach analyzes a fixed set of logs from real driving scenes, which provides insights into the system’s performance across naturally occurring scenarios and their distributions.
Figure 1: Overview of the proposed validation process. We embed a scenario using a pretrained encoder $\mathcal{E}$. The embedding is concatenated with a difficulty score from a predictor head $\mathcal{D}$ which takes the embedding as input. Concatenated embeddings from a set of scenarios are clustered and weighted. Importance sampling and simulation is performed to compute validation metrics on autonomy (red dots are selected scenarios).
@@
  • The sampling method should return the exposure to each scenario type to calculate a weighted validation metric.
@@ II Related Work
AV Validation & Verification. ISO 26262 is a functional safety standard for automotive electronic and electrical systems, focusing on lifecycle management and risk assessment. ISO 26262 sets the limit for the acceptable ratio of faults as 10 Failures in Time (FIT), meaning that given $10^{9}$ hours of operation no more than 10 faults should be observed. However, AVs are also generally unsafe to test on the road at that scale and modern AVs undergo extremely dynamic and fast-paced engineering cycles that necessitate frequent re-verification and validation [68, 69]. Simulation has emerged as a partial solution for AV V&V that is rapid, scalable, and safe compared to real-world testing. Plenty of open-source simulators exist for AV testing [70, 71, 72, 73, 74] with some including benchmarks for autonomous driving quality [75, 76]. Despite being more scalable than real-world testing, simulated driving can still be prohibitively expensive at scale. A standard approach is to mine for difficult scenarios from existing driving logs [1, 2, 3, 4], or synthesize difficult or rare scenarios often using ML [77, 78, 79, 80, 81, 82, 83, 84] with [85] showing off a controllable diffusion-based approach allowing users to specify desired scenario properties. Fault injection is another common approach for creating difficult test scenarios, with frameworks such as [86, 87, 88, 89, 90, 91, 92] using algorithms like Bayesian Optimization to rapidly hunt for faults that lead to system failures. Other approaches specifically target ML-based components by inserting sensor noise [93, 94], with “white-box” frameworks like [95, 96] adversarially mutating inputs in an attempt to maximize “neuron coverage” in the system to be tested. Formal verification methods typically attempt to give rigorous guarantees on system safety by using mathematical tools like Lyapunov functions [97, 98, 99] and Satisfiability Modulo Theories [100, 101], or logical modelling and proofs [102, 103, 104, 105, 106].
@@ III-A1 Tracks
The track tensor $X_{T}\in\mathbb{R}^{N_{T}\times T\times D_{T}}$ contains $N_{T}$ tracks from the perception system, including classes such as vehicles, pedestrians, cones, lane dividers, and others. For each track-time step, we encode the pose, velocity, acceleration, extents, and class into a single $D_{T}$-dimensional vector. We assign each unique track to a row in $X_{T}$. If there are fewer than $N_{T}$ tracks or a track is not visible at all time steps, $X_{T}$ is zero-padded where necessary.
@@ III-A2 Traffic Signals
The signal tensor $X_{S}\in\mathbb{R}^{N_{S}\times T\times D_{S}}$ contains $N_{S}$ signals where each time step has a $D_{S}$-vector encoding pose and a perception label for each signal. We use a map with signal locations, so no time steps are padded as they can be set with the pose and an UNKNOWN label.
@@ III-B Time Invariant Inputs
To encode the road network, we adopt an approach similar to [5]. We extract vectorized polylines representing the centers of lanes and annotation marks such as stop lines, parking spaces, and crosswalks. The $N_{Z}$ polylines are represented with three tensors:
  • Coordinate frames for each polyline $X_{F}\in\mathbb{R}^{N_{Z}\times D_{F}}$ where each $D_{F}$-vector contains $x,y$ locations and sine and cosine of the rotation to the frame
  • Class labels $X_{L}\in\mathbb{R}^{N_{Z}\times D_{L}}$ for one-hot encoded $D_{L}$-vector
  • Sets of $S_{Z}$ polyline points $X_{P}\in\mathbb{R}^{N_{Z}\times S_{Z}\times D_{P}}$ expressed in the associated coordinate frame where each $D_{P}$-vector contains $x,y$ positions, width, and existence
@@ III-C Masking
As is typical in an MAE setup, we randomly sample inputs to mask with masking ratio $r$. Masking corresponds to replacing the encoded features’ vectors with a 0-vector. For the time variant masks $M_{T}\in\mathbb{R}^{N_{T}\times T}$ and $M_{S}\in\mathbb{R}^{N_{S}\times T}$, we sample across the first two dimensions of the input independently. For the time invariant mask $M_{Z}\in\mathbb{R}^{N_{Z}}$, we randomly mask over the polylines. If a polyline is masked, the coordinate frame is left unmasked but the labels and points are all masked, including the existence. Note that at inference time, we do not mask the inputs.
@@ III-D Model
  • The inputs are not tokenized nor of varying sizes, so we project them into a shared space and back out to their original shape
@@ III-E Loss
We independently sample a mask with ratio $r_{\text{loss}}$ following the same sampling methodology as described previously in the masking section. For all inputs that are covered by this mask, we compute a reconstruction loss. Thus, the loss applies to both masked and unmasked inputs. The loss is a weighted sum with weights $\lambda$ over individual input types:
@@
In addition to the track loss $\mathcal{L}_{T}$, signal loss $\mathcal{L}_{S}$, and road network loss $\mathcal{L}_{R}$, we introduce a loss term $\mathcal{L}_{\text{ego}}$ for reconstructing ego, as tasks may require a strong representation of ego. Each loss consists of an L1 loss on $J$ continuous inputs $x_{j}$, and a cross-entropy loss on $M$ categorical inputs $x_{m}$:
@@
Where $\hat{x}_{j}$ is a predicted continuous input and $\hat{x}_{m}$ is a predicted categorical input. For the models used in later sections, we set all $\lambda=1$.
@@ IV-A Data
The data for the fine-tuning is similar to data used to train the difficulty model presented in [111]. We use a set of recorded driving logs and their results after simulation. For a given scenario, we apply a label of 1 if simulation resulted in a collision and 0 if it did not. In general, it is possible for a driving log to have a simulation result on many different software versions. If so, we add it to the dataset as multiple examples. Note that each would share identical inputs, as we encode the original recorded driving log. For scenarios that do not unimodally fail or succeed, the optimal score is then some real number between 0 and 1.
@@ IV-B Fine-tuning
While log divergence and other factors can lead to false positive collisions in our training set, the simulator is of sufficient quality where many simulated collisions are worth measuring.
@@ V-A Clustering scenarios
During validation, it is crucial to ensure that every kind of scenario, especially those less common or more complex, is tested to validate changes in behavior comprehensively. After clustering scenarios, we can expose the autonomy stack to a representative sample from each cluster to cover each of the various behaviors on which validation is required. This intuition motivates a sampling technique where we uniformly sample scenarios across clusters (see Algorithm 1).
@@ V-B Importance sampling
We incorporate scenario difficulty (Section IV) by appending and weighting the scenario’s difficulty score as part of the embedded space. Thus, clustering will separate based on scenario similarity and difficulty.
With clusters separated in terms of difficulty, we can score each cluster (e.g. according to an average, or even a particular percentile of the difficulty scores within the cluster). The score $w_{i}$ for cluster $c_{i}$ can then be treated as an importance weight upon sampling. The resulting sampling scheme, DICE, presented in Algorithm 2, ensures that the autonomy stack is more frequently exposed to challenging scenarios, which are crucial for thorough validation.
@@ VII Conclusions and Limitations
To illustrate how expensive validation can be, consider a hypothetical example validating that a driving stack violates a safety metric (measured in miles per violation) no more than once per $N$ miles using a real-time simulator. If all roads have a 35 mph speed limit and we wish to simulate 10 times the desired metric for confidence, this requires $10N/35\approx 0.286N$ driving hours of simulation. At time of writing, the on-demand cost for an AWS g5.4xlarge node is \$$\$1.624$ per hour [113] and we assume one node fits one concurrent simulation. This yields a cost of \$$\$0.464N$. For a violation that should occur fewer than once every $50{,}000$ miles, this would cost roughly \$$\$23{,}200$ for each code change.
There are many nuances to effective large scale simulation that are not included here, such as validating separately in different regions of a map, running multiple simulations per node, faster than real time simulation, node startup and shutdown costs, data transmission costs, limitations on pipeline width, and simulation fidelity.